OpenStudy (anonymous):

series expansion question: find the power series expansion for the function x/(2-x) and find the radius of convergence.

OpenStudy (anonymous):

I know the power series formula is 1/1-x and my study guide says to get there factor out the 1/2 in x/(2-x)....which is ok but the next step in my guide just disappears the x on top like this: x/2−x=(1/2)∗1/(1−(x/2)) howwhywhat happened to the x in the numerator?

OpenStudy (amistre64):

i would do long division to put this into a power series

OpenStudy (amistre64):

x/2 +x^2/4+x^3/8 .... ------------------ 2-x ) x (x-x^2/2) ----------- x^2/2 (x^2/2 -x^3/4) --------------- x^3/4

OpenStudy (amistre64):

as such we notice a pattern \[\sum_{n=1}^{inf}\frac{x^{n}}{2^n}\]

OpenStudy (anonymous):

heh, YOU notice a pattern =-)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sum+x%5En%2F2%5En+%2C+n%3D1+to+inf urgh .... i got another negative lol

OpenStudy (anonymous):

not again...lol

OpenStudy (anonymous):

we have singular point is 2 we need center point?

OpenStudy (amistre64):

using this summation we can find the rest of it

OpenStudy (amistre64):

\[\lim_{n\to\ inf}\frac{-x^n}{2^n}\frac{2^{n-1}}{-x^{n-1}}\]

OpenStudy (amistre64):

the negatives cancel and it looks like all our ns disapear to make this a 1

OpenStudy (amistre64):

\[|\frac{x}{2}|<1\] \[|x|<2\]

OpenStudy (amistre64):

as such, our radius is 2 and our interval of convergence is -2 to 2

OpenStudy (anonymous):

If you think of the x^n/2^n as (x/2)^n absolute value of x/2 converges when less than 1 and its less than 1 when x < 2 as amistre64 says

OpenStudy (amistre64):

lol, one of these days i mmight even figure out what im doing :)

OpenStudy (anonymous):

we have singular point is 2 if power series expansion at 0 ->radius is 2

OpenStudy (amistre64):

sorry chanh, i got no way to verify your method :/

OpenStudy (amistre64):

at least not in generality

OpenStudy (anonymous):

and the disappearing n's (and the corresponding -x) explains why the x's disappeared in my study key they just skipped like one billion steps

OpenStudy (amistre64):

skipping steps saves on ink

OpenStudy (anonymous):

I'm thinking of just drawing a black box on my test labelled magic - and putting the answer next to it.

OpenStudy (amistre64):

the thunder was a nice touch, if you can recreate that effect during the test, your sure to get an A lol

OpenStudy (anonymous):

I'm study in complex variables with applications book

OpenStudy (anonymous):

hmmm I dont know how that lines up with Early Transcendentals